OpenStudy (anonymous):

one angle of a triangle is 15 degrees greater than the smallest angle, and the third is 35 degrees less than twice the smallest. find the measurements smallest- ? second angle- ? third angle ?

OpenStudy (lgbasallote):

since the smallest angle was mentioned in both...let x = smallest one angle = 15 + x third angle = 2x - 35 these three must equal 180 so... x + 15 + x + 2x - 35 = 180 ill let you do the rest :D

OpenStudy (campbell_st):

draw a picture  then using angle sum of a triangle 2x - 35 + x+15 + x = 180 4x - 20 = 180 I'll let you solve the equation which will give the size of the smallest angle... you'll need to work out the other 2... just look at the diagram